# Improving Medium Range Severe Weather Prediction through Transformer Post-processing of AI Weather Forecasts

### v2

@@ 5 Case Studies
The top 10 largest positive and negative BSSs are provided in Table 3, along with the number of 80-km grid boxes containing $\geq$ 1 storm report for each event. We present two cases here from the Top 5 BSSs: 17 April 2024 (BSS of 0.12; Fig. 9) and 1 May 2024 (BSS of -0.16; Fig. 10). The 17 April 2024 event had the second highest aggregate BSS of all cases and is an example where DOT(Pangu-GFS) provided excellent medium-range guidance for a spatially compact severe weather event. The DOT(Pangu-GFS) forecasts were superior to the DOT(GFS) forecasts for the Day 4–5 lead-times (c.f., Figs. 9b,f and Fig. 9c,g), capturing the focused area of severe weather reports across MI, OH, and PA up to 5 days in advance and providing consistent guidance between Day 3–5. At lead-times of 4–6 days, the maximum hazard forecast probabilities within the GFS_no_convect forecast were shifted the south (Fig. 9f–h), for the Day 4–6 forecasts, eventually focusing on the correct area by Day 3 (Fig. 9e).
BSS $>$ 0 BSS $<$ 0
@@
The 1 May 2024 event had the largest negative BSS, indicating the most significant difference in aggregate Brier Scores between all pairs of forecasts in favor of the DOT(GFS) forecasts. In this case, DOT(Pangu-GFS) produced large areas of false alarms over KS and MO, with forecast probabilities shifted too far to the east relative to the corridor of severe weather reports across western KS, OK, and TX (Fig. 10a–d). This is especially notable over west TX, where the Day 4–6 DOT(Pangu-GFS) forecasts had probabilities $<$ 5% where reports occurred (Fig. 10b–d). In this event, severe convection was initiated by a dryline (not shown), a feature that may not have been adequately represented in the underlying Pangu-GFS forecasts compared to the GFS. By Day 3, the differences between the forecasts were less pronounced, although an eastward bias was still present in DOT(Pangu-GFS) (c.f., Fig. 10a,e).
@@ 6 Summary and Discussion
Further, combining multiple models into an ensemble produced the most robust forecasts, with the operational model mean consistently achieving the highest skill across all metrics and forecast days. This highlights the potential for integrating diverse model outputs to enhance operational forecast performance and mitigate individual model weaknesses. Feature attribution analysis using Integrated Gradients showed that even without explicit convective parameters, Pangu-Weather was able to leverage information from standard meteorological variables to effectively predict severe weather across multiple forecast days.
Case studies revealed both strengths and limitations of AI-based forecasts. DOT(Pangu-GFS) excelled at predicting spatially compact severe weather events several days in advance, it occasionally produced eastward-biased forecasts and struggled with certain meteorological phenomena like drylines. Additionally, reliability analysis showed that individual models tend to slightly over-forecast at higher probability thresholds, though this tendency diminishes at longer lead times.
Future work should explore incorporating a wider variety of AI models, developing methods to derive convective parameters directly from AI model outputs, and further investigating the meteorological factors driving model performance differences. The framework presented here offers a promising pathway for operational implementation of AI-enhanced medium-range severe weather prediction, potentially improving both forecast accuracy and lead time for high-impact events.
@@ Acknowledgements
NWS Storm Reports that were used for model training were retrieved from the NOAA SPC storm report archive located at https://www.spc.noaa.gov/wcm/#data. The Pangu-Weather model is available at https://github.com/198808xc/Pangu-Weather. Initial conditions, including ERA5, HRES, and GFS analyses were retrieved from the NCAR Research Data Archive located at https://rda.ucar.edu.

### v3

@@ 5 Case Studies
The top 10 largest positive and negative BSSs are provided in Table 3, along with the number of 80-km grid boxes containing $\geq$ 1 storm report for each event. We present two cases here from the Top 5 BSSs: 17 April 2024 (BSS of 0.12; Fig. 10) and 1 May 2024 (BSS of -0.16; Fig. 11). The 17 April 2024 event had the second highest aggregate BSS of all cases and is an example where DOT(Pangu-GFS) provided excellent medium-range guidance for a spatially compact severe weather event. The DOT(Pangu-GFS) forecasts were superior to the DOT(GFS) forecasts for the Day 4–5 lead-times (c.f., Figs. 10b,f and Fig. 10c,g), capturing the focused area of severe weather reports across MI, OH, and PA up to 5 days in advance and providing consistent guidance between Day 3–5. At lead-times of 4–6 days, the maximum hazard forecast probabilities within the DOT(GFS) forecast were shifted the south (Fig. 10f–h), for the Day 4–6 forecasts, eventually focusing on the correct area by Day 3 (Fig. 10e).
BSS >> 0 BSS << 0
@@
The 1 May 2024 event had the largest negative BSS, indicating the most significant difference in aggregate Brier Scores between all pairs of forecasts in favor of the DOT(GFS) forecasts. In this case, DOT(Pangu-GFS) produced large areas of false alarms over KS and MO, with forecast probabilities shifted too far to the east relative to the corridor of severe weather reports across western KS, OK, and TX (Fig. 11a–d). This is especially notable over west TX, where the Day 4–6 DOT(Pangu-GFS) forecasts had probabilities << 5% where reports occurred (Fig. 11b–d). In this event, severe convection was initiated by a dryline (not shown), a feature that may not have been adequately represented in the underlying Pangu-GFS forecasts compared to the GFS. By Day 3, the differences between the forecasts were less pronounced, although an eastward bias was still present in DOT(Pangu-GFS) (c.f., Fig. 11a,e).
@@ 6 Summary and Discussion
Further, combining multiple models into an ensemble produced the most robust forecasts, with the operational model mean consistently achieving the highest skill across all metrics and forecast days. This highlights the potential for integrating diverse model outputs to enhance operational forecast performance and mitigate individual model weaknesses. Feature attribution analysis using Integrated Gradients showed that even without explicit convective parameters, Pangu-Weather was able to leverage information from standard meteorological variables to effectively predict severe weather across multiple forecast days. However, limitation of our current attribution visualization and analysis is the temporal aggregation across 24-hour periods, which may obscure important subdaily variations in feature importance. Future investigations should consider implementing subdaily attribution analysis, similar to the tree interpreter approach of Mazurek et al. (2025), to better understand the temporal evolution of feature contributions and validate the meteorological reasonableness of observed attribution patterns.
Case studies revealed both strengths and limitations of AI-based forecasts. DOT(Pangu-GFS) excelled at predicting spatially compact severe weather events several days in advance, it occasionally produced eastward-biased forecasts and struggled with certain meteorological phenomena like drylines. Additionally, reliability analysis showed that individual models tend to slightly over-forecast at higher probability thresholds, though this tendency diminishes at longer lead times. While the AI models demonstrated promising skill during the unusually active 2024 season, a longer verification period spanning multiple years is necessary to fully assess the generalizability of these results and quantify skill variations across different climatological regimes.
Future work should explore incorporating a wider variety of AI models, developing methods to derive convective parameters directly from AI model outputs, and further investigating the meteorological factors or non meteorological factors like elevation variability Hua and Chavas (2019) and vegetation Kellner and Niyogi (2014) driving model performance differences. Understanding how errors and biases in the underlying AI models propagate through post-processing is key to improving these systems. Nevertheless, the framework presented here offers a promising pathway for operational implementation of AI-enhanced medium-range severe weather prediction, potentially improving both forecast accuracy and lead time for high-impact events.
@@ Acknowledgements
NWS Storm Reports that were used for model training were retrieved from the NOAA SPC storm report archive located at https://www.spc.noaa.gov/wcm/#data. The Pangu-Weather model is available at https://github.com/198808xc/Pangu-Weather. Initial conditions, including ERA5, HRES, and GFS analyses were retrieved from the NCAR Research Data Archive located at https://rda.ucar.edu. The code implementation and model checkpoints will be available on the Zenodo repository upon acceptance of this work.
@@
Figure S1: Similar to Fig. 3 but for experiment using DOT(Pangu-HRES) by either reducing “d_model” from the 256 in the study down to 128, 64 and 32 or reducing “num_layers” from 4 down to 3,2 and 1. No significance difference in performance with respect to the reference configuration in the formal analysis.
Figure S2: Similar to Fig. 7 but use DOT(GFS+Convect) as comparison baseline